OpenStudy (anonymous):

I know that any number except 0 is equal to 1 when multiplied to the 0 power. What happens when 0^0?

Parth (parthkohli):

The same thing that happens when you do 0/0.

Parth (parthkohli):

I'm serious about that answer though. There are multiple ways to think about 0^0, such as the following:\[0^0\]\[= 0^{(1-1)}\]\[= 0^1 \div 0^1\]\[= 0 \div 0\]I'm not gonna write the next step, or the mathies are gonna put me into prison.